CBC Montreal broadcasts 4 hours of live news from Monday to Friday and 2 hours per day Saturday and Sunday. You are tasked to assess the storage hardware cost of archiving the news. Assume that the programs are encoded HD at a rate of 20 Mbps. A 24 TB (Tera-byte) storage unit (Drobo 5N2 24TB) sells on Amazon for $2,760. What would be the annual cost of hardware if you use this unit (5 Marks).
Answer #1

Cost of 24TB = $2,760

20 Mb is stored per second.

There are 52 weeks and 1 day in a year

Assuming the extra day is Sunday or Saturday

Channel takes 52*(4*5+2*2)+1*2 for a week=1250 hrs = 1250 * 60 * 60 = 4500000 seconds

Total data stored = 20 * 4500000 = 90000000 Mb

Number of 24 TB Hard disk used = 90000000Mb/24Tb= 720/24 = 30 disks

Therefore, annual cost of hard-disks is 30 * 2760 = $82800

If the extra day is Monday to Friday

Channel takes 52*(4*5+2*2)+1*4 for a week=1252 hrs = 1252 * 60 * 60 = 4507200 seconds

Total data stored = 20 * 4507200 = 90144000 Mb

Number of 24 TB Hard disk used = 90144000Mb/24Tb=721.152/24 = 30.048 disks

So consider 31 disks to store data

Therefore, annual cost of hard-disks is 31 * 2760 = $85560